FOURIER TRANSFORM INFRARED SPECTRAL INVESTIGATION OF THE ν_{6} BAND OF CYCLIC-C_{3}H_{2}

\underline{PRADEEP RISIKRISHNA VARADWAJ}, RYUJI FUJIMORI, KENTAROU KAWAGUCHI, Department of Chemistry, Faculty of Science, Okayama University, 3-1-1, Tsushima-Naka, Okayama 700-8530, JAPAN.

The gas phase absorption spectrum of the ν_{6} band (out of plane in phase CH bend) of cyclopropenylidene (c-C_{3}H_{2}) has been observed using a high-resolution Fourier transform infrared spectrometer for the first time. The molecule was produced by microwave discharge through a flow of allene (25 mTorr) and Ar (30 mTorr) mixture inside a glass cell. The observed spectrum shows c-type ro-vibrational transitions in which the Q-branch lines (J values up to 34 and K_{a} values up to 8) are strongly and distinctly stand out in the spectrum. A least squares fitting of a total of 357 tranitions (332 ro-vibrational transitions from this work and 25 transitions from the millimeter-wave spectrum with 10^{6} times larger statistical weight) results in the determination of the rotational and centrifugal distortion constants in the upper state. The rotational constants agree with those determined by millimter-wave spectrum^{a} and thereby confirming the ν_{6} band (ν_{0} = 776.11729(17) cm^{−1}) of c-C_{3}H_{2}.

aR. Mollaaghababa, C. Gottlieb, J. M. Vrtilek, and P. E. Thaddeus, J. Chem. Phys. 99, 890 (1993)